# Highest-weight vectors and three-point functions in GKO coset decomposition

### v1

@@ 5 Kyiv formula for Painlevé $\mathrm{III}_{3}$ tau-function
In this section we deduce Kyiv formulas for the tau function from the coset (or blowup) relations (4.59) following [Nek24], [JN20]. We restrict ourselves to the case of Painlevé $\mathrm{III}_{3}$ which corresponds to Whittaker limit of conformal blocks. First, we recall Hamiltonian of the Painlevé $\mathrm{III}_{3}$ and show the relation between tau function and generating function of canonical transformation. Then we define Whittaker vectors and Whittaker limits of conformal blocks. The Whittaker limit of $\widehat{\mathfrak{sl}}(2)$ conformal block satisfies the non-stationary affine Toda equation. In the classical $k\rightarrow\infty$ this leads to solution of Hamilton-Jacobi equation for Painlevé $\mathrm{III}_{3}$ Hamiltonian. Taking the classical limit of relations (4.59) we get the Kyiv formula.
@@ 5.1 Generating function and tau function for Painlevé $\mathrm{III}_{3}$
Consider extended phase space with coordinates $(x,p,z)$ where $x$ is coordinate,$p$ momentum, $z$ time with degenerate Poisson bracket defined by $\{x,p\}=1,\{x,z\}=\{p,z\}=0$. Hamiltonian dynamics is defined by one function which is called Hamiltonian $H(x,p;z)$ and differential equations called Hamilton equations
@@
Assume not that there is another pair of functions $\alpha,\beta$ on the extended phase space and function $S(x,\alpha;z)$ such that
@@
and Hamilton-Jacobi equation
@@
Here and below by additional indices in partial derivatives, we emphasize the variables that are considered to be fixed. The function $S(x,\alpha;z)$ is a called a generating function of a canonical transformation from coordinates $(x,p,z)$ to $(\alpha,\beta,z)$. The following result is standard
@@
The Hamiltonain of the Painlevé $\mathrm{III}_{3}$ equation has the form (see e.g. [GIL13])
@@
Suppose $S(x,\alpha;z)$ satisfy Hamilton-Jacobi equation (5.1) for Painlevé $\mathrm{III}_{3}$, $\beta,p$ are defined via (5.2) and $\alpha,\beta$ are integrals of motion. Then
$\log\tau=S-\alpha\beta-4z\left(\frac{\partial S}{\partial z}\right)_{\alpha,x}-2x\left(\frac{\partial S}{\partial x}\right)_{\alpha,z}$ (5.6)
@@ 5.2 Whittaker limit of three-point functions
(a) If $\lambda,k$ are generic then there exists a unique up to constant $\widehat{\mathfrak{sl}}(2)$ Whittaker vector, moreover $\langle\mathcal{W}(x,z)|v_{\lambda,k}\rangle\neq 0$.
(b) If $P,b$ are generic then there exists a unique up to constant Virasoro Whittaker vector, moreover $\langle\mathbb{W}(z)|v_{P,b}\rangle\neq 0$.
@@
$\Psi_{k}(\lambda;x,z)=\langle\mathcal{W}_{\lambda}^{k}(1,1),\mathcal{W}_{\lambda}^{k}(x,z)\rangle.$ (5.12)
@@
$\mathrm{F}_{b}(P;z)=\langle\mathbb{W}_{P}^{b}(1),\mathbb{W}_{P}^{b}(z)\rangle.$ (5.13)
@@
$\langle\mathbb{W}_{l}(z)\otimes\mathcal{W}_{l}(x,z),\tilde{u}_{l}(\lambda)\rangle=\|\tilde{u}_{l}(\lambda)\|^{2}z^{\Delta(P(\lambda),b)+l^{2}}x^{-l-\lambda/2}.$ (5.15)
@@ 5.3 Kyiv formula
$z\partial_{z}\Psi_{k}(\lambda;x,z)=\frac{1}{\kappa}\Big{(}zx^{-1}+x+x^{2}\partial_{x}^{2}\Big{)}\Psi_{k}(\lambda;x,z)$ (5.17)
Of course, this is just a Knizhnik-Zamolodchikov equation in the Whittaker limit
@@
where $S(\sigma;x,z),S_{0}(\sigma;x,z)$ are certain functions that does not depend on $\kappa$. The following proposition follows directly from the Toda equation (5.17),
@@
$\tau(\sigma,\beta;z)=\sum_{l\in\mathbb{Z}}\frac{1}{\mathtt{t}^{1,0}_{2l}(-2\sigma)^{2}}e^{l\beta}\mathrm{F}_{1}(i(\sigma+l);z).$ (5.22)
The formula (5.22) is called the Kyiv formula, it was conjectured in [GIL12], [GIL13]. There are several other proofs this formula, namely [ILT15], [BS15], [GL18a], [GL18b]; as was mentioned above here we follow the logic of [Nek24], [JN20].
@@
Therefore the leading divergent contributions of the left side and right side are equal to $\exp\Big{(}-\kappa S(\sigma;x,z)\Big{)}$ and cancels. In the subleading order we get
$e^{(1-\sigma\partial_{\sigma}-2x\partial_{x}-4z\partial_{z})S(\sigma,x;z)}=\sum_{l\in\mathbb{Z}}\frac{1}{\mathtt{t}^{1,0}_{2l}(-2\sigma)^{2}}e^{-l\partial_{\sigma}S(\sigma;x,z)}\mathrm{F}_{1}(i(\sigma+l);z).$ (5.24)
Let $\beta=-\partial_{\sigma}S(\sigma,x;z)$, $p=\partial_{x}S(\sigma,x;z)$. Now we can apply Theorem 5.1. Namely we know that $S$ satisfies Hamilton-Jacobi equation and assume that $\sigma,\beta$ are integrals of motion, then $x.p$ satisfies dynamics (5.1) with Painlevé $\mathrm{III}_{3}$ Hamiltonian. Furthermore, by Proposition 5.3 the left side is equal to the tau function. Putting all things together we get formula (5.22). ∎
@@ 6.1 Integral representation of vertex operators
$\mathcal{V}_{\nu}(x,z)=\int\limits_{0\leq t_{N}\leq\dots\leq t_{1}\leq z}\!\!\!\!(1+x\gamma(z))^{\nu}:\!e^{\frac{\nu}{\sqrt{2\kappa}}\varphi(z)}\!:\prod_{i=1}^{N}S(t_{i})dt_{i},$ (6.1)
@@
$\mathcal{U}_{\nu,n}(x,z)=:\!e^{-g_{0}}e^{n\sqrt{2}\phi(z)}(-1)^{nh_{0}^{(1)}(h_{0}^{(1)}-1)}e^{g_{0}}\!:\sum_{r\geq 0}\binom{\nu+2n}{r}x^{r}\mathcal{O}^{(N)}_{r;\nu}(z).$ (6.2)
@@
$\mathcal{U}_{\nu,n}(x,z)=\int\limits_{0\leq t_{N}\leq\dots\leq t_{1}\leq z}\!\!\!\!\!\!\!\!\Big{(}e^{-g_{0}}e^{n\sqrt{2}\phi(z)}(-1)^{nh_{0}^{(1)}(h_{0}^{(1)}-1)}e^{g_{0}}\Big{)}(1+x\gamma(z))^{\nu+2n}:\!e^{\frac{\nu}{\sqrt{2\kappa}}\varphi(z)}\!:\prod_{i=1}^{N}S(t_{i})dt_{i}.$ (6.3)
@@
$\mathcal{U}_{\nu,-{1}/{2}}(x,z)=\Big{(}b_{1}(z)\mathcal{V}_{\nu}(x,z)-\frac{1}{\nu}(b_{0}(z)+xb_{1}(z))\partial_{x}\mathcal{V}_{\nu}(x,z)\Big{)}\\ =\int\limits_{0\leq t_{N}\leq\dots\leq t_{1}\leq z}\!\!\!\!:\!\Big{(}b_{1}(z)-xb_{0}(z)\gamma(z)\Big{)}(1+x\gamma(z))^{\nu-1}e^{\frac{\nu}{\sqrt{2\kappa}}\varphi(z)}\!:\prod_{i=1}^{N}S(t_{i})dt_{i}\\ =\int\limits_{0\leq t_{N}\leq\dots\leq t_{1}\leq z}\!\!\!\!:\!\Big{(}e^{-g_{0}}b_{1}(z)e^{g_{0}}\Big{)}(1+x\gamma(z))^{\nu-1}e^{\frac{\nu}{\sqrt{2\kappa}}\varphi(z)}\!:\prod_{i=1}^{N}S(t_{i})dt_{i}.$ (6.4)
@@
The Theorem 6.1 can be viewed as an operator analog of Theorem 3.7. We had two proofs of the latter, one based on explicit computation and another based on the operator $I(z)$. Similarly, one can expect two proofs of the Theorem 6.1.
@@
$e^{g_{0}}e^{\Delta}(w)e^{-g_{0}}=\beta(w),\quad e^{g_{0}}f^{\Delta}(w)e^{-g_{0}}=f^{\Delta}(w)+h^{(1)}(w)\gamma(w)+\gamma^{\prime}(w).$ (6.6)
@@
$e^{g_{0}}\mathit{U}_{\nu,n}(x,z)e^{-g_{0}}=\mathcal{V}_{\nu,n}(x,z)=:\!e^{n\sqrt{2}\phi(z)}(-1)^{nh_{0}^{(1)}(h_{0}^{(1)}-1)}\!:\sum_{p\geq 0}\binom{\nu+2n}{p}x^{p}\mathcal{O}^{(0)}_{p;\nu}(z).$ (6.7)
For the commutation with $\beta_{r}$ only $\mathcal{O}^{(0)}_{r;\nu}(z)$ are important. Hence using (2.53) we have
@@
It remains to compute (using (2.53))
@@
$:\!P^{(1)}\big{[}\partial\phi\big{]}e^{m\sqrt{2}\phi(z)}(-1)^{mh_{0}^{(1)}(h_{0}^{(1)}-1)}\!:\otimes:\!P^{(2)}\big{[}\partial\varphi,\beta,\gamma\big{]}e^{\frac{\nu}{\sqrt{2\kappa}}\varphi(z)}\!:$ (6.12)
where $P^{(1)}$, $P^{(2)}$ are two differential polynomials (c.f. formula (2.42)). The space of fields for arbitrary $x$ is obtained by conjugation by $e^{xf_{0}^{\Delta}}$. But due to relation (6.5c) we have
$\mathit{U}_{\nu,n}(x,z)=e^{xf_{0}^{\Delta}}\mathit{U}_{\nu,n}(0,z)e^{-xf_{0}^{\Delta}},$ (6.13)
@@
Step 3. The space of fields can be identified with the space of states $L_{i}\otimes\mathcal{M}_{\nu}$. We know that $\mathit{U}_{\nu,n}(x,z)$ is a field, in Step 1 we showed that the corresponding vector is the highest weight vector for $\mathfrak{sl}(2)_{k+1}$ with highest weight $\nu+2n$. Moreover, it is straightforward to see that the corresponding vector has a conformal dimension equal to the conformal dimension of $u_{n}(\nu)$ There exists only one up to proportionally vector with this property in $L_{2\{n\}}\otimes\mathcal{M}_{\nu}$. It remains to check the normalization. ∎
@@
$\mathcal{U}_{1,-{1}/{2}}(x,z)=\Big{(}e^{-g_{0}}b_{1}(z)e^{g_{0}}\Big{)}e^{\frac{1}{\sqrt{2\kappa}}\varphi(z)}=I(z)$ (6.14)
@@
$\mathcal{U}_{1,-\frac{1}{2}}(x,z)\mathcal{U}_{\nu,n}(y,w)\\ =e^{-g_{0}}b_{1}(z)e^{g_{0}}e^{\frac{1}{\sqrt{2\kappa}}\varphi(z)}\Big{(}e^{-g_{0}}e^{n\sqrt{2}\phi(w)}(-1)^{nh_{0}^{(1)}(h_{0}^{(1)}-1)}e^{g_{0}}\Big{)}(1+y\gamma(w))^{\nu+2n}:\!e^{\frac{\nu}{\sqrt{2\kappa}}\varphi(w)}\!:\\ =\pm(z-w)^{d(n,\nu,k)}\Big{(}\Big{(}e^{-g_{0}}e^{(n-\frac{1}{2})\sqrt{2}\phi(w)}(-1)^{(n-\frac{1}{2})h_{0}^{(1)}(h_{0}^{(1)}-1)}e^{g_{0}}\Big{)}(1+y\gamma(w))^{\nu+2n}:\!e^{\frac{\nu+1}{\sqrt{2\kappa}}\varphi(w)}\!:\\ +O\left(z-w\right)\Big{)}$ (6.18)

### v3

@@ 5 Kyiv formula for Painlevé $\mathrm{III}_{3}$ tau-function
In this section we deduce Kyiv formulas for the tau function from the coset (or blowup) relations (4.59) closely following [Nek24], [JN20]. We restrict ourselves to the case of Painlevé $\mathrm{III}_{3}$ which corresponds to Whittaker limit of conformal blocks. First, we recall Hamiltonian of the Painlevé $\mathrm{III}_{3}$ and show the relation between tau function and generating function of canonical transformation. Then we define Whittaker vectors and Whittaker limits of conformal blocks. The Whittaker limit of $\widehat{\mathfrak{sl}}(2)$ conformal block satisfies the non-stationary affine Toda equation. In the classical $k\rightarrow\infty$ this leads to solution of Hamilton-Jacobi equation for Painlevé $\mathrm{III}_{3}$ Hamiltonian. Taking the classical limit of relations (4.59) we get the Kyiv formula.
@@ 5.1 Generating function and tau function for Painlevé $\mathrm{III}_{3}$
Consider extended phase space with coordinates $(x,p,z)$ where $x$ is a coordinate, $p$ is a momentum, and $z$ is a time. The degenerate Poisson bracket is defined by $\{x,p\}=1,\{x,z\}=\{p,z\}=0$. Hamiltonian dynamics is defined by one function which is called Hamiltonian $H(x,p;z)$ and differential equations called Hamilton equations
@@
Here and below by additional indices in partial derivatives, we emphasize the variables that are considered to be fixed.
Assume that there is another pair of functions $\alpha,\beta$ on the extended phase space and a function $S(x,\alpha;z)$ such that
@@
and Hamilton-Jacobi equation holds
@@
The function $S(x,\alpha;z)$ is a called a generating function of a canonical transformation from coordinates $(x,p,z)$ to $(\alpha,\beta,z)$. The following result is standard
@@
The Hamiltonian of the Painlevé $\mathrm{III}_{3}$ equation has the form (see e.g. [GIL13])
@@
Suppose function $S(x,\alpha;z)$ satisfies Hamilton-Jacobi equation (5.1) for Painlevé $\mathrm{III}_{3}$ Hamiltonian, functions $\beta,p$ are defined by relations (5.2) and functions $\alpha,\beta$ are integrals of motion. Then
$\log\tau=S-\alpha\beta-4z\left(\frac{\partial S}{\partial z}\right)_{\alpha,x}-2x\left(\frac{\partial S}{\partial x}\right)_{\alpha,z}.$ (5.6)
@@ 5.2 Whittaker limit of three-point functions
(a) If $\lambda,k$ are generic then there exists a unique up to constant $\widehat{\mathfrak{sl}}(2)$ Whittaker vector, moreover $\big{\langle}\mathcal{W}_{\lambda}^{k}(x,z),v_{\lambda,k}\big{\rangle}\neq 0$.
(b) If $P,b$ are generic then there exists a unique up to constant Virasoro Whittaker vector, moreover $\big{\langle}\mathbb{W}_{P}^{b}(z),v_{P,b}\big{\rangle}\neq 0$.
@@
$\Psi_{k}(\lambda;x,z)=\Big{\langle}\mathcal{W}_{\lambda}^{k}(1,1),\mathcal{W}_{\lambda}^{k}(x,z)\Big{\rangle}.$ (5.12)
@@
$\mathrm{F}_{b}(P;z)=\Big{\langle}\mathbb{W}_{P}^{b}(1),\mathbb{W}_{P}^{b}(z)\Big{\rangle}.$ (5.13)
@@
$\langle\mathbb{W}_{l}(z)\otimes\mathcal{W}_{l}(x,z),\tilde{u}_{l}(\lambda)\rangle=\|\tilde{u}_{l}(\lambda)\|^{2}z^{\Delta(P(\lambda),b)+l^{2}}x^{-l-\lambda/2},$ (5.15)
@@ 5.3 Kyiv formula
$z\partial_{z}\Psi_{k}(\lambda;x,z)=\frac{1}{\kappa}\Big{(}zx^{-1}+x+x^{2}\partial_{x}^{2}\Big{)}\Psi_{k}(\lambda;x,z).$ (5.17)
Of course, this is just a Knizhnik-Zamolodchikov equation in the Whittaker limit.
@@
where $S(\sigma;x,z),S_{0}(\sigma;x,z)$ are certain functions that do not depend on $\kappa$. The following proposition follows directly from the Toda equation (5.17).
@@
$\tau(\sigma,\beta;z)=\sum_{l\in\mathbb{Z}}\frac{1}{\mathtt{t}^{1,0}_{2l}(-2\sigma)^{2}}{\mathrm{e}}^{l\beta}\mathrm{F}_{1}({\mathrm{i}}(\sigma+l);z).$ (5.22)
The formula (5.22) is called the Kyiv formula, it was first conjectured in [GIL12], [GIL13]. There are several other proofs this formula, namely [ILT15], [BS15], [GL18a], [GL18b]; as was mentioned above here we follow the logic of [Nek24], [JN20].
@@
Therefore the leading divergent contributions of the left side and right side are equal to $\exp\Big{(}-\kappa S(\sigma;x,z)\Big{)}$ and cancel. In the subleading order we get
${\mathrm{e}}^{(1-\sigma\partial_{\sigma}-2x\partial_{x}-4z\partial_{z})S(\sigma,x;z)}=\sum_{l\in\mathbb{Z}}\frac{1}{\mathtt{t}^{1,0}_{2l}(-2\sigma)^{2}}{\mathrm{e}}^{-l\partial_{\sigma}S(\sigma;x,z)}\mathrm{F}_{1}({\mathrm{i}}(\sigma+l);z).$ (5.24)
Let $\beta=-\partial_{\sigma}S(\sigma,x;z)$, $p=\partial_{x}S(\sigma,x;z)$. Now we can apply Theorem 5.1. Namely we know that $S$ satisfies Hamilton-Jacobi equation and assume that $\sigma,\beta$ are integrals of motion, then $x,p$ satisfy dynamics (5.1) with Painlevé $\mathrm{III}_{3}$ Hamiltonian. Furthermore, by Proposition 5.3 the left side is equal to the tau function. Putting all things together we get formula (5.22). ∎
@@ 6.1 Integral representation of vertex operators
$\mathcal{V}_{\nu}(x,z)=\int\limits_{0\leq t_{N}\leq\dots\leq t_{1}\leq z}\!\!\!\!(1+x\gamma(z))^{\nu}:\!{\mathrm{e}}^{\frac{\nu}{\sqrt{2\kappa}}\varphi(z)}\!:\prod_{i=1}^{N}S(t_{i})dt_{i},$ (6.1)
@@
$\mathcal{U}_{\nu,n}(x,z)={\mathrm{e}}^{-g_{0}}:\!{\mathrm{e}}^{n\sqrt{2}\phi(z)}\!\!:(-1)^{nh_{0}^{(1)}(h_{0}^{(1)}-1)}{\mathrm{e}}^{g_{0}}\sum_{r\geq 0}\binom{\nu+2n}{r}x^{r}\mathcal{O}^{(N)}_{r;\nu}(z).$ (6.2)
@@
$\mathcal{U}_{\nu,n}(x,z)=\!\!\!\int\limits_{0\leq t_{N}\leq\dots\leq t_{1}\leq z}\!\!\!\!\!\!\!\!\Big{(}{\mathrm{e}}^{-g_{0}}:\!{\mathrm{e}}^{n\sqrt{2}\phi(z)}\!\!:(-1)^{nh_{0}^{(1)}(h_{0}^{(1)}-1)}{\mathrm{e}}^{g_{0}}\Big{)}(1+x\gamma(z))^{\nu+2n}:\!{\mathrm{e}}^{\frac{\nu}{\sqrt{2\kappa}}\varphi(z)}\!\!:\prod_{i=1}^{N}S(t_{i})dt_{i}.$ (6.3)
@@
$\mathcal{U}_{\nu,-{1}/{2}}(x,z)=\Big{(}b_{1}(z)\mathcal{V}_{\nu}(x,z)-\frac{1}{\nu}(b_{0}(z)+xb_{1}(z))\partial_{x}\mathcal{V}_{\nu}(x,z)\Big{)}\\ =\int\limits_{0\leq t_{N}\leq\dots\leq t_{1}\leq z}\!\!\!\!\Big{(}b_{1}(z)-b_{0}(z)\gamma(z)\Big{)}(1+x\gamma(z))^{\nu-1}:\!{\mathrm{e}}^{\frac{\nu}{\sqrt{2\kappa}}\varphi(z)}\!\!:\prod_{i=1}^{N}S(t_{i})dt_{i}\\ =\int\limits_{0\leq t_{N}\leq\dots\leq t_{1}\leq z}\!\!\!\!\Big{(}{\mathrm{e}}^{-g_{0}}b_{1}(z){\mathrm{e}}^{g_{0}}\Big{)}(1+x\gamma(z))^{\nu-1}:\!{\mathrm{e}}^{\frac{\nu}{\sqrt{2\kappa}}\varphi(z)}\!\!:\prod_{i=1}^{N}S(t_{i})dt_{i}.$ (6.4)
@@
Theorem 6.1 can be viewed as an operator analog of Theorem 3.7. We had two proofs of the latter, one based on explicit computation and another one based on the operator $I(z)$. Similarly, one can expect two proofs of the Theorem 6.1.
@@
${\mathrm{e}}^{g_{0}}e^{\Delta}(w){\mathrm{e}}^{-g_{0}}=\beta(w),\quad{\mathrm{e}}^{g_{0}}f^{\Delta}(w){\mathrm{e}}^{-g_{0}}=f^{\Delta}(w)+h^{(1)}(w)\gamma(w)+\gamma^{\prime}(w).$ (6.6)
@@
${\mathrm{e}}^{g_{0}}\mathit{U}_{\nu,n}(x,z){\mathrm{e}}^{-g_{0}}=\mathcal{V}_{\nu,n}(x,z)=:\!{\mathrm{e}}^{n\sqrt{2}\phi(z)}\!\!:(-1)^{nh_{0}^{(1)}(h_{0}^{(1)}-1)}\sum_{p\geq 0}\binom{\nu+2n}{p}x^{p}\mathcal{O}^{(0)}_{p;\nu}(z).$ (6.7)
For the commutation with $\beta_{r}$ only $\mathcal{O}^{(0)}_{p;\nu}(z)$ are important. Hence using (2.53a) we have
@@
It remains to compute (using (2.53c))
@@
$:\!P^{(1)}\big{[}\partial\phi\big{]}{\mathrm{e}}^{m\sqrt{2}\phi(z)}(-1)^{mh_{0}^{(1)}(h_{0}^{(1)}-1)}\!:\otimes:\!P^{(2)}\big{[}\partial\varphi,\beta,\gamma\big{]}{\mathrm{e}}^{\frac{\nu}{\sqrt{2\kappa}}\varphi(z)}\!:$ (6.12)
where $P^{(1)}$, $P^{(2)}$ are two differential polynomials (c.f. formula (2.42)). The space of fields for arbitrary $x$ is obtained by conjugation by ${\mathrm{e}}^{xf_{0}^{\Delta}}$. But due to relation (6.5c) we have
$\mathit{U}_{\nu,n}(x,z)={\mathrm{e}}^{xf_{0}^{\Delta}}\mathit{U}_{\nu,n}(0,z){\mathrm{e}}^{-xf_{0}^{\Delta}},$ (6.13)
@@
Step 3. The space of fields can be identified with the space of states $L_{i}\otimes\mathcal{M}_{\nu}$. We know that $\mathit{U}_{\nu,n}(x,z)$ is a field, in Step 1 we showed that the corresponding vector is the highest weight vector for $\mathfrak{sl}(2)_{k+1}$ with highest weight $\nu+2n$. Moreover, it is straightforward to see that the corresponding vector has a conformal dimension equal to the conformal dimension of $u_{n}(\nu)$. There exists only one up to proportionally vector with this property in $L_{2\{n\}}\otimes\mathcal{M}_{\nu}$. It remains to check the normalization. ∎
@@
$\mathcal{U}_{1,-{1}/{2}}(x,z)=\Big{(}{\mathrm{e}}^{-g_{0}}b_{1}(z){\mathrm{e}}^{g_{0}}\Big{)}{\mathrm{e}}^{\frac{1}{\sqrt{2\kappa}}\varphi(z)}=I(z)$ (6.14)
@@
$\mathcal{U}_{1,-\frac{1}{2}}(x,z)\mathcal{U}_{\nu,n}(y,w)\\ ={\mathrm{e}}^{-g_{0}}b_{1}(z){\mathrm{e}}^{g_{0}}:\!{\mathrm{e}}^{\frac{1}{\sqrt{2\kappa}}\varphi(z)}\!\!:\Big{(}{\mathrm{e}}^{-g_{0}}:\!{\mathrm{e}}^{n\sqrt{2}\phi(w)}\!\!:(-1)^{nh_{0}^{(1)}(h_{0}^{(1)}-1)}{\mathrm{e}}^{g_{0}}\Big{)}(1+y\gamma(w))^{\nu+2n}:\!{\mathrm{e}}^{\frac{\nu}{\sqrt{2\kappa}}\varphi(w)}\!\!:\\ =\pm(z-w)^{d(n,\nu,k)}\Big{(}\Big{(}{\mathrm{e}}^{-g_{0}}:\!{\mathrm{e}}^{(n-\frac{1}{2})\sqrt{2}\phi(w)}\!\!:(-1)^{(n-\frac{1}{2})h_{0}^{(1)}(h_{0}^{(1)}-1)}{\mathrm{e}}^{g_{0}}\Big{)}(1+y\gamma(w))^{\nu+2n}:\!{\mathrm{e}}^{\frac{\nu+1}{\sqrt{2\kappa}}\varphi(w)}\!\!:\\ +O\left(z-w\right)\Big{)}$ (6.18)